### v2

@@ 4.1 Experimental Setup
Datasets   We perform extensive experiments across three reasoning tasks and six datasets: (1) general visual question answering: ChartQA Masry et al. (2022) and MMStar Chen et al. (2024b); (2) mathematical reasoning: MathVista Lu et al. (2023), MathVerse Zhang et al. (2025), and MathVision Wang et al. (2024a); (3) commonsense and scientific reasoning: GAOKAO-MM Zong & Qiu (2024).
Models   To demonstrate the versatility of AStar, we evaluate its effectiveness on both LLM and MLLM, including Qwen2.5-7B Qwen Team (2024), Qwen2-VL-2B Wang et al. (2024b), and Qwen2-VL-7B. This design aims to validate that AStar can seamlessly leverage pre-trained LLM and MLLM as its inference backbone without modifications.
Baselines   We evaluate AStar against four strong baseline categories: (1) open-source general MLLMs, including the powerful Qwen2-VL Wang et al. (2024b) and InternVL2 series Chen et al. (2024d); (2) open-source MLLMs specifically optimized for mathematical reasoning, such as recent work URSA Luo et al. (2025) and Math-LLaVA Shi et al. (2024); (3) advanced closed-source MLLMs, including GPT-4V OpenAI (2023), and GPT-4o OpenAI (2024); and (4) tree-based methods, such as AR-MCTS Dong et al. (2024a) and Mulberry Yao et al. (2024).
Table 1: Evaluation of AStar’s reasoning capabilities on MathVista and MathVerse testmini. The best results are highlighted in bold. For MathVista, we pick five categories: ALL (overall accuracy), ARI (arithmetic reasoning), LOG (logical reasoning), STA (statistical reasoning), and VQA (visual question answering). For MathVerse, we present seven categories: ALL (overall accuracy), VI (vision intensive), VD (vision dominant), VO (vision only), TD (text dominant), TL (text lite), and TO (text only).
Model #Params MathVista MathVerse
ALL ARI LOG STA VQA ALL VI VD VO TD TL TO
Random - 17.9 13.8 13.4 14.3 26.3 12.4 12.4 12.4 12.4 12.4 12.4 12.4
Human - 60.3 59.2 40.7 63.9 55.9 64.9 61.4 68.3 66.7 71.2 70.9 41.7
@@
mPLUG-Owl2-7B Ye et al. (2023) 7B 22.2 19.2 13.5 21.4 27.9 10.3 11.1 9.4 8.0 11.6 11.4 13.8
MiniGPT4-7B Zhu et al. (2023) 7B 23.1 32.0 10.8 17.9 30.2 12.2 12.5 14.8 8.7 12.3 12.9 13.4
LLaVA-1.5-13B Liu et al. (2024a) 13B 27.7 28.6 10.8 22.9 30.2 12.7 12.6 12.7 9.0 17.1 12.0 22.6
SPHINX-V2-13B Lin et al. (2023) 13B 36.7 33.4 24.3 51.5 43.0 16.1 16.4 15.6 16.2 20.8 14.1 14.0
SPHINX-MoE Lin et al. (2023) 8$\times$7B 42.6 43.0 14.4 50.8 43.3 22.8 21.1 19.6 18.3 33.3 21.9 23.1
LLaVA-NeXT-34B Liu et al. (2024b) 34B 46.5 - - - - 34.6 35.2 28.9 22.4 49.0 37.6 30.1
InternLM-XComposer2-VL Dong et al. (2024c) 7B 57.6 51.6 13.5 62.8 39.7 25.9 20.1 24.4 19.8 36.9 28.3 42.5
Deepseek-VL Lu et al. (2024) 7B 34.9 38.8 18.9 33.2 34.6 19.3 20.2 18.4 11.8 23.0 23.2 23.1
InternVL2-8B Chen et al. (2024d) 8B 58.3 56.4 10.8 68.8 49.7 35.9 32.2 30.9 27.7 39.0 33.8 36.0
Qwen2-VL Wang et al. (2024b) 7B 58.9 57.5 24.3 43.1 58.1 33.6 31.3 30.3 28.1 37.4 33.5 35.0
@@
G-LLaVA-7B Gao et al. (2023) 7B 25.1 19.4 15.2 15.1 28.7 16.6 17.2 14.6 9.4 20.9 20.7 21.1
Math-LLaVA-13B Shi et al. (2024) 13B 46.6 40.7 23.3 42.3 33.5 22.9 24.5 21.7 16.1 27.3 24.9 27.0
Math-PUMA-Qwen2-7B Zhuang et al. (2024) 7B 47.9 46.2 21.6 55.8 30.2 33.6 33.4 31.6 26.0 42.1 35.0 39.8
Math-PUMA-DeepSeek-Math Zhuang et al. (2024) 7B 44.7 41.9 8.1 50.8 31.3 31.8 33.6 31.6 14.7 43.4 35.4 47.5
MAVIS-7B Zhang et al. (2024d) 7B - - - - - 35.2 34.1 29.7 31.8 43.2 37.2 -
InfiMM-Math Han et al. (2024) 7B - - - - - 34.5 38.1 32.4 15.8 46.7 32.4 -
MultiMath-7B Peng et al. (2024) 7B 50.0 42.2 23.3 64.9 49.2 27.7 28.1 25.9 15.0 34.8 30.8 35.3
URSA-7B Luo et al. (2025) 7B 59.8 53.5 21.6 57.1 40.2 45.7 46.4 43.9 28.6 55.3 48.3 51.8
AStar (Ours, Training-free Reasoning) 7B 63.5 63.1 61.3 68.8 60.1 54.0 46.8 61.8 46.4 53.9 44.3 53.9
Table 2: Comparison with powerful baselines on general VQA datasets: MMStar and ChartQA.
Model #Params MMStar ChartQA
2B-Scale Baselines
Qwen2-VL Wang et al. (2024b) 2B 48.0 73.5
Mulberry Yao et al. (2024) 2B 51.3 77.7
AStar (Ours) 2B 51.7 78.3
$\geq$ 7B-Scale Baselines
Deepseek-VL Lu et al. (2024) 7B 36.1 59.1
Qwen2-VL Wang et al. (2024b) 7B 60.7 83.0
InternVL2 Chen et al. (2024d) 8B 61.5 83.3
Insight-V Dong et al. (2024d) 7B 61.5 82.3
Mulberry Yao et al. (2024) 7B 61.3 83.9
LLaVA-CoT Xu et al. (2024) 11B 58.1 -
LlamaV-o1 Thawakar et al. (2025) 11B 59.5 -
AStar (Ours) 7B 61.7 83.9
@@ 4.2 Performance Comparison with Baselines
Results on Diverse Reasoning Benchmarks   Table 1 and 2 presents the performance of AStar across four mainstream reasoning benchmarks. We have four key findings:
$\diamondsuit$ AStar consistently outperforms both powerful open-source general and math-specialized MLLMs across general visual question answering and mathematical reasoning tasks. Specifically, using Qwen2.5-7B as the reasoning backbone, our method achieves $54.0\%$ accuracy on MathVerse, surpassing InternVL2-8B by $18.1\%$ and the recent high-quality CoT-enhanced URSA-7B model by $8.3\%$.
$\diamondsuit$ AStar demonstrates strong performance across multiple reasoning capabilities. Notably, it achieves $63.1\%$ accuracy on logical reasoning (LOG), outperforming InternVL2-8B by $50.5\%$ and URSA-7B by $39.7\%$. Similar improvements are observed in statistical reasoning (STA: $68.8\%$) and visual question answering (VQA: $60.1\%$).
@@
$\diamondsuit$ Beyond specialized reasoning tasks, AStar performs superior on general VQA benchmarks (MMStar and ChartQA), suggesting that enhanced reasoning capabilities also benefit general visual understanding.
Table 3: Comparison with leading multimodal tree search methods. AStar achieves competitive performance with only 0.5K prior data and 5 search iterations per sample, demonstrating superior efficiency and effectiveness.
Methods Open-Source Only Training-Free Data Volume Search Iter. $\downarrow$ MathVista Acc. $\uparrow$ MMStar Acc. $\uparrow$ GAOKAO Acc. $\uparrow$
AR-MCTS Dong et al. (2024a) 34.5K 32.0 64.1 - 37.4
Mulberry Yao et al. (2024) 260K - 63.1 61.3 -
@@
[…]
Figure 4: Comparison with leading MLLMs across various capabilities on the challenging MathVision dataset. Our 7B-parameter model achieves comparable performance to GPT-4o.
@@ 4.3 Data and Computational Efficiency
We further compare our approach AStar with two recent powerful multimodal tree search methods: AR-MCTS Dong et al. (2024a) and Mulberry Yao et al. (2024). As shown in Table 3, we compare multiple dimensions, including training requirements, training data volume, computational efficiency (measured by average search iterations per sample), and accuracy on three representative benchmarks.
Results demonstrate that AStar achieves competitive performance while significantly improving both data and computational efficiency. In terms of data efficiency, AStar requires only 0.5K prior samples, which represents a 69-fold reduction compared to AR-MCTS (34.5K) and a 520-fold reduction compared to Mulberry (260K), respectively. This efficiency stems from explicit reasoning pattern extraction, enabling performance comparable to implicit learning methods with far less training data. Such efficiency is particularly advantageous in limited-data scenarios. Regarding computational efficiency, AStar reduces the average iterations per sample by 6.4×6.4$\times$6.4 × compared to AR-MCTS, while maintaining comparable or superior accuracy across all benchmarks.
Figure 5: Comparison between AStar and powerful MLLMs across 3 challenging benchmarks: MathVista, MathVerse, and MathVision. ‘OS’ and ‘CS’ denote open-source and closed-source models. AStar with 7B model outperforms larger OS and CS models.
Figure 6: Cross-distribution performance comparison on GAOKAO-MM. We test on Qwen2-VL-2B, Qwen2-VL-7B, and GPT-4o. Results highlight that AStar exhibits superior performance when testing on out-of-distribution data.
@@ 4.4 Out-of-Distribution Generalization
Recent work has highlighted that distributional shifts severely affect the reliability of MLLMs Yin et al. (2023); Zhang et al. (2024f). While these models excel in in-distribution (ID) tasks, their performance often degrades in out-of-distribution (OOD) scenarios Dong et al. (2024b); Miyai et al. (2024), a challenge further compounded by the difficulty of acquiring sufficient high-quality training data.
@@ 4.5 Ablation Study and Analysis
  To investigate the contribution of each component in AStar, we conduct ablation study as shown in Table 4. Here, “w/o” denotes variants with specific components removed. Our analysis reveals three key findings:
Table 4: Ablation results on AStar-7B. ‘RA’, ‘RC’, ‘RS’, ‘SC’ denotes ‘random actions’, ‘random card’, ‘random selection’, and ‘self-consistency’, respectively. We observe that every component is important for optimal performance.
Model Setting MathVista MathVerse Average
AStar 63.5 54.0 58.8
   -- w/o thought cards (RA) 58.9 48.5 53.7-5.1-${}^{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{\textbf{-5.1}\downarrow}}$
   -- w/o card match (RC) 61.3 50.3 55.8-3.0-${}^{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{\textbf{-3.0}\downarrow}}$
   -- w/o verification (RS) 62.0 51.6 56.8-2.0-${}^{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{\textbf{-2.0}\downarrow}}$
   -- w/o verification (SC) 63.0 51.8 57.4-1.4-${}^{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{\textbf{-1.4}\downarrow}}$
@@
Inference Time Scaling   We gradually increase the number of selected thought cards during inference to investigate whether our method follows test-time scaling laws. Figure 7 demonstrates that incorporating additional reasoning guidance via thought cards leads to consistent performance improvements in our AStar framework.
Figure 7: Inference Time Scaling. We examine the variation in AStar performance with the number of selected reasoning guidance-providing thought cards in Sec. 3.3.
@@
Visualization of Sampling Space   To analyze the diversity of solutions generated by our AStar-powered reasoning framework across varying numbers of thought cards, we conduct a visualization study here. We sample 250 problems from MathVision and generate 2 candidate solutions for each problem using AStar (Qwen2-VL-2B), resulting in 500 total samples. Following Dong et al. (2024a), we employ a three-step analysis process using BGE-M3 (Chen et al., 2024a) for semantic embedding, followed by dimensionality reduction (PCA) and clustering techniques, DBSCAN Ester et al. (1996).
@@
Figure 8 and 9 illustrate the visualization results across varying numbers of thought cards ($k=2$ and $k=5$) on MathVista and MathVision, respectively. The empirical analysis reveals consistent conclusions across both standard mathematical reasoning tasks (MathVista) and challenging scenarios (MathVision). When fewer thought cards are provided ($k=2$), the model generates clusters with fewer centroids for the same problem set, indicating a more constrained sampling space with limited potential. Conversely, increasing the number of thought cards to $k=5$ leads to clusters with more centroids, suggesting a richer sampling space with higher performance potential. The modest difference in cluster counts between $k=2$ and $k=5$ indicates that our framework effectively adapts to problem complexity by dynamically selecting suitable reasoning patterns from high-quality thought cards. Even with fewer thought cards, it maintains robust sampling diversity and performance through this adaptive matching mechanism.
@@
Figure 8: Visualization of candidate reasoning paths on MathVista with $k=2$ (left) and $k=5$ (right) thought cards. Dots represent individual reasoning steps, while colors indicate distinct solution clusters. The increasing density and diversity of clusters from left to right demonstrate the model’s enhanced exploration capability with more thought cards, leading to more comprehensive coverage of the solution space.
@@
Figure 9: Visualization of candidate reasoning paths on MathVision with $k=2$ (left) and $k=5$ (right) thought cards. Dots represent individual reasoning steps, while colors indicate distinct solution clusters. The increasing density and diversity of clusters from left to right demonstrate the model’s enhanced exploration capability with more thought cards, leading to more comprehensive coverage of the solution space.
Notably, the observed lower cluster count in MathVista compared to MathVision offers valuable insights into task-specific reasoning requirements. On MathVista, which contains relatively simpler problems, our AStar framework exhibits strong reasoning path convergence: multiple high-quality thought cards typically lead to consistent solutions. This convergence allows for effective performance even with basic self-consistency verification. However, on the more challenging MathVision dataset, we observe greater path divergence, emphasizing the importance of careful validation and selection of optimal reasoning paths. This behavior closely resembles human problem-solving patterns. When addressing elementary problems, different approaches tend to naturally converge to the same answer. In contrast, for complex challenges like Olympic problems, diverse approaches often yield different conclusions, making thorough solution validation necessary.
In summary, our visualization results empirically validate that AStar effectively addresses the challenge of limited diversity in candidate solutions, enabling comprehensive coverage of the problem-solving space while providing robust prior guidance for efficient inference.

### v3

@@ 3.1 Experimental Setup
Datasets   We perform extensive experiments across three tasks and eight datasets: (1) mathematical reasoning: MathVista [36], MathVerse [37], and MathVision [5]; (2) commonsense and scientific reasoning: GAOKAO-MM [38]; (3) general visual perception and understanding: ChartQA [39], MMStar [40], MMMU [41] and BLINK [42]. For all benchmarks, we adopt the official evaluation metrics. Further details can be found in Appendix D.1.
Models   To demonstrate AStar’s versatility, we evaluate its effectiveness on both LLM and MLLM, including Qwen2.5-7B [43], and Qwen2-VL-2/7B [44]. This design aims to validate that AStar can seamlessly leverage pre-trained LLM and MLLM as its inference backbone without modifications.
Baselines   In this paper, we comprehensively evaluate AStar against four strong and representative baseline categories: (1) open-source general MLLMs, including the powerful Qwen2-VL [45] and InternVL2 series [46]; (2) open-source MLLMs specifically optimized for complex reasoning, such as SFT-based methods URSA [13], Math-LLaVA [47], and RLVR-based methods R1-VL-7B [18], LMM-R1 [20], and MM-Eureka [48]; (3) explicit search-based methods such as Mulberry [15]; and (4) advanced closed-source MLLMs, including GPT-4o [2], and GPT-4V [49].
Table 1: Evaluation of AStar’s reasoning abilities on MathVista and MathVerse testmini. The best results are highlighted in bold. For MathVerse, we show 7 categories: ALL (overall accuracy), VI (vision intensive), VD (vision dominant), VO (vision only), TD (text dominant), TL (text lite), and TO (text only). For MathVista, we present 5 categories: ALL (overall accuracy), ARI (arithmetic reasoning), LOG (logical reasoning), STA (statistical reasoning), and VQA (visual question answering).
Model #Params MathVerse MathVista
ALL VI VD VO TD TL TO ALL ARI LOG STA VQA
Random - 12.4 12.4 12.4 12.4 12.4 12.4 12.4 17.9 13.8 13.4 14.3 26.3
Human - 64.9 61.4 68.3 66.7 71.2 70.9 41.7 60.3 59.2 40.7 63.9 55.9
@@
MiniGPT4-7B [50] 7B 12.2 12.5 14.8 8.7 12.3 12.9 13.4 23.1 32.0 10.8 17.9 30.2
LLaVA-1.5-13B [51] 13B 12.7 12.6 12.7 9.0 17.1 12.0 22.6 27.7 28.6 10.8 22.9 30.2
SPHINX-V2-13B [52] 13B 16.1 16.4 15.6 16.2 20.8 14.1 14.0 36.7 33.4 24.3 51.5 43.0
SPHINX-MoE [52] 8$\times$7B 22.8 21.1 19.6 18.3 33.3 21.9 23.1 42.6 43.0 14.4 50.8 43.3
LLaVA-NeXT-34B [53] 34B 34.6 35.2 28.9 22.4 49.0 37.6 30.1 46.5 - - - -
InternLM-XComposer2-VL [54] 7B 25.9 20.1 24.4 19.8 36.9 28.3 42.5 57.6 51.6 13.5 62.8 39.7
Deepseek-VL [55] 7B 19.3 20.2 18.4 11.8 23.0 23.2 23.1 34.9 38.8 18.9 33.2 34.6
InternVL2-8B [46] 8B 35.9 32.2 30.9 27.7 39.0 33.8 36.0 58.3 56.4 10.8 68.8 49.7
Qwen2-VL [45] 7B 33.6 31.3 30.3 28.1 37.4 33.5 35.0 58.9 57.5 24.3 43.1 58.1
@@
G-LLaVA-7B [56] 7B 16.6 17.2 14.6 9.4 20.9 20.7 21.1 25.1 19.4 15.2 15.1 28.7
Math-LLaVA-13B [47] 13B 22.9 24.5 21.7 16.1 27.3 24.9 27.0 46.6 40.7 23.3 42.3 33.5
Math-PUMA-Qwen2-7B [57] 7B 33.6 33.4 31.6 26.0 42.1 35.0 39.8 47.9 46.2 21.6 55.8 30.2
Math-PUMA-DeepSeek-Math [57] 7B 31.8 33.6 31.6 14.7 43.4 35.4 47.5 44.7 41.9 8.1 50.8 31.3
MAVIS-7B [58] 7B 35.2 34.1 29.7 31.8 43.2 37.2 - - - - - -
InfiMM-Math [59] 7B 34.5 38.1 32.4 15.8 46.7 32.4 - - - - - -
MultiMath-7B [60] 7B 27.7 28.1 25.9 15.0 34.8 30.8 35.3 50.0 42.2 23.3 64.9 49.2
URSA-8B [13] 7B 45.7 46.4 43.9 28.6 55.3 48.3 51.8 59.8 53.5 21.6 57.1 40.2
R1-VL-7B [18] 7B 40.0 37.3 33.6 39.8 45.0 40.1 40.7 63.5 54.7 28.5 61.2 60.9
AStar (Qwen2.5-7B, Training-free) 7B 53.9 49.7 64.4 48.6 56.4 48.3 56.1 64.2 63.8 59.5 69.1 60.9
AStar (Qwen2-VL-7B, Training-free) 7B 47.5 41.8 51.6 48.6 49.3 51.0 42.2 61.7 60.6 28.6 68.4 59.6
@@ 3.2 Performance Comparison with Baselines
Foundational Reasoning Enhancement Analysis   Table 1 presents the performance of AStar across two mainstream reasoning benchmarks. We have three key findings:
$\circ$ AStar consistently outperforms both general and math-specialized MLLMs. With Qwen2.5-7B reasoning backbone, AStar achieves $53.9\%$ accuracy on MathVerse, surpassing large-scale CoT-trained URSA-7B by $8.2\%$, and the computationally-intensive GRPO-trained R1-VL-7B by $13.9\%$.
$\circ$ AStar demonstrates strong performance across diverse reasoning capabilities. It reaches $59.5\%$ on logical reasoning (LOG), outperforming the powerful URSA-7B by $37.9\%$. Similar gains are observed in statistical reasoning (STA: $69.1\%$) and visual question answering (VQA: $60.9\%$).
@@
Figure 3: Results on the challenging reasoning benchmark, MathVision. AStar-7B (Qwen2.5-7b) achieves competitive performance to GPT-4o.
Model Perception Comprehension Understanding MMStar BLINK MMMU ChartQA Qwen2-VL-2B 48.0 44.4 41.1 73.5 AStar (Ours, 2B) 52.0 46.7 42.4 78.6 Qwen2-VL-7B 60.7 53.2 54.1 83.0 AStar (Ours, 7B) 61.7 56.0 57.2 83.9
Table 2: Evaluation results on visual perception and understanding benchmarks: (1) Visual Perception: MMStar and BLINK; (2) Context Comprehension: MMMU; (3) Chart Understanding: ChartQA. Our framework generalizes beyond multimodal reasoning to visual perception and understanding domains.
Table 3: Comparison with two primary methods: search-based methods (Mulberry) and post-training methods (LMM-R1, MM-Eureka, R1-VL, URSA). ‘OS Only’ indicates datasets constructed only with open-source models. As a training-free reasoning framework, AStar achieves superior performance with only 500 prior samples, demonstrating superior efficiency and effectiveness.
Methods Type OS Only Training-Free Prior Data $\downarrow$ MMStar Acc. $\uparrow$ MathVerse Acc. $\uparrow$ MathVision Acc. $\uparrow$
Mulberry-7B [15] Search 260K 61.3 - -
URSA-8B [13] SFT 1100K 55.4 45.7 26.2
LMM-R1-3B [20] GRPO 55.3K 58.0 41.6 26.4
MM-Eureka-7B [48] GRPO 15K 59.4 50.3 26.9
R1-VL-7B [18] GRPO 260K 60.0 40.0 27.1
Ours (Qwen2.5-7B) - 0.5K 62.3 53.9 32.7
Ours (Qwen2-VL-7B) - 0.5K 61.7 47.5 26.8
@@
Figure 4: Comparison between AStar and powerful MLLMs across 3 challenging benchmarks: MathVista, MathVerse, and MathVision. ‘OS’ and ‘CS’ denote open-source and closed-source models. AStar with 7B backbone outperforms them.
Extending Reasoning Gains to Visual Perception and Understanding   To investigate whether our thought-augmented paradigm can generalize to visual perception and understanding tasks, we conduct experiments on four representative benchmarks in Table 2, including visual perception (MMStar [40] and BLINK [42]), contextual comprehension (MMMU [41]), and chart understanding (ChartQA [39]). Results show that our reasoning framework significantly enhances performance across these domains and models of different scales. This reveals that thought-driven structured thinking mechanisms not only benefit complex problem-solving in multimodal reasoning but also strengthen fundamental visual understanding abilities even in less explicitly reasoning-intensive tasks.
@@ 3.3 Data and Computational Efficiency
We further compare AStar with two primary categories of reasoning methods: search-based (Mulberry) and post-training approaches (URSA, LMM-R1, MM-Eureka, R1-VL). Table 3 evaluates these methods across training requirements, open-source compatibility, training data volume, and performance on two tasks: visual understanding (MMStar) and reasoning (MathVerse, MathVision).
[…]
Figure 5: Cross-distribution performance comparison on GAOKAO-MM. We test on Qwen2-VL-2B, Qwen2-VL-7B, and GPT-4o. Results highlight AStar’s superior out-of-distribution performance.
@@ 3.4 Out-of-Distribution Generalization
Recent work has highlighted that distributional shifts severely affect the reliability of MLLMs [29, 61]. While these models excel in in-distribution (ID) tasks, their performance often degrades in out-of-distribution (OOD) scenarios [62]. This challenge is compounded by the inherent difficulty of acquiring sufficient training data and computational resources.
@@ 3.5 Ablation Study and Analysis
Ablation Study   To investigate the contribution of each component in AStar, we conduct ablation study as shown in Table 4. Here, “w/o” denotes variants with specific components removed. Our analysis reveals three key findings:
@@
Model Setting MathVista MathVerse Average AStar 64.2 53.9 59.1 -- w/o thought cards (RAC) 59.4 47.9 53.6-5.5-${}^{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{\textbf{\scriptsize-5.5}\downarrow}}$- w/o card match (RC) 61.7 49.8 55.8-3.3-${}^{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{\textbf{\scriptsize-3.3}\downarrow}}$- w verification (RS) 62.8 51.7 57.2-1.9-${}^{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{\textbf{\scriptsize-1.9}\downarrow}}$- w verification (SC) 63.7 52.0 57.9-1.2-${}^{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{\textbf{\scriptsize-1.2}\downarrow}}$
Table 4: Ablation results on AStar-7B (Qwen2.5-7B). ‘RAC’, ‘RC’, ‘RS’, ‘SC’ denotes ‘random action combinations’, ‘random card’, ‘random selection’, and ‘self-consistency’, respectively. We observe that every component is important for optimal performance.
Figure 6: Inference time scaling results.
@@
Inference Time Scaling   We gradually increase the number of selected thought cards during inference to investigate whether our method follows test-time scaling laws. Figure 6 shows that incorporating more reasoning guidelines via thought cards leads to consistent performance improvements.
Figure 7: Results on Qwen2.5-VL-3B. AStar integration with RLVR-trained models further enhances performance.
Plug-and-Play Capability   Given that AStar operates as a test-time inference framework, it can be effectively combined with post-training methods. We applied our reasoning paradigm to models that had already undergone post-training procedures (RLVR-trained LMM-R1 [20], i.e., Reinforcement Learning with Verifiable Rewards). As shown in Figure 7, our framework not only seamlessly integrates with these approaches but also consistently enhances their performance. This synergistic effect indicates that AStar could capture unique reasoning patterns complementary to those learned during post-training, further highlighting AStar’s plug-and-play nature. Future work could explore deeper integration with post-training methods to maximize complementary benefits.
@@
Visualization of Sampling Space To analyze the diversity of solutions generated by our AStar-powered reasoning framework across varying numbers of thought cards, we conduct a visualization study here. We sample 250 problems from MathVision and generate 2 candidate solutions for each problem using AStar (Qwen2-VL-2B), resulting in 500 total samples. Following prior work [14], we employ a three-step analysis process using BGE-M3 [63] for semantic embedding, followed by dimensionality reduction and clustering techniques.
Step 1: Semantic Embedding:   We utilize the BGE-M3 encoder $f_{e}$ to map each answer $a_{i}$ from our solution set $A={a_{1},...,a_{n}}$ into a high-dimensional semantic space, producing embedding vectors $\mathbf{e}_{i}\in\mathbb{R}^{d}$:
$\mathbf{e}_{i}=f_{e}(a_{i}),\quad i\in{1,...,n}$ (9)
where $n=500$ denotes our total sample size, and $d$ represents the dimension of the embedding space.
Step 2: Dimensionality Reduction:   We implement a two-stage dimensionality reduction process, beginning with z-score normalization:
$\hat{\mathbf{e}}_{i}=\frac{\mathbf{e}_{i}-\mu}{\sigma}$ (10)
where $\mu$ and $\sigma$ represent the mean and standard deviation vectors across all embeddings. We then apply PCA to project the standardized embeddings into a two-dimensional space:
$\mathbf{v}_{i}=\text{PCA}(\hat{\mathbf{e}}_{i})\in\mathbb{R}^{2}$ (11)
Step 3: Density-Based Clustering:   To identify distinct solution patterns, we apply DBSCAN [64] to the reduced embeddings:
$C_{k}=\text{DBSCAN}({\mathbf{v}i}{i=1}^{n};\epsilon,\text{minPts})$ (12)
where $\epsilon=0.7$ defines the neighborhood distance threshold and $\text{minPts}=4$ specifies the minimum points required for a dense region. $C_{k}$ denotes the cluster set identified for the number of thought cards $k$.
@@
Figure 8 and 9 illustrate the visualization results across varying numbers of thought cards ($k=2$ and $k=5$) on MathVista and MathVision datasets, respectively. The empirical analysis reveals consistent conclusions across both standard mathematical reasoning tasks (MathVista) and challenging scenarios (MathVision). When fewer thought cards are provided ($k=2$), the model generates clusters with fewer centroids for the same problem set, indicating a more constrained sampling space with limited potential. Conversely, increasing the number of thought cards to $k=5$ leads to clusters with more centroids, suggesting a richer sampling space with higher performance potential.
@@
Figure 8: The visualization of candidate reasoning paths across different numbers of thought cards on MathVista. Dots represent individual reasoning steps, while colors indicate distinct solution clusters. The comparison shows results for $k=2$ (left) and $k=5$ (right) thought cards, respectively. The increasing density and diversity of clusters from left to right demonstrate the model’s enhanced exploration capability with more thought cards, leading to more comprehensive coverage of the solution space.
@@
Figure 9: The visualization of candidate reasoning paths across different numbers of thought cards on MathVision. Dots represent individual reasoning steps, while colors indicate distinct solution clusters. The comparison shows results for $k=2$ (left) and $k=5$ (right) thought cards, respectively. The increasing density and diversity of clusters from left to right demonstrate the model’s enhanced exploration capability with more thought cards, leading to more comprehensive coverage of the solution space.
Notably, the difference in cluster count between $k=2$ and $k=5$ is not particularly large. This observation validates that our framework can effectively adapts to problem complexity by dynamically selecting the most suitable reasoning patterns from high-quality thought cards. Even with fewer thought cards, the framework maintains robust sampling diversity and performance through this adaptive matching mechanism.
Other Analysis   Finally, we include some additional results and analysis in the Appendix, including a more comprehensive performance comparison (E.1, E.2), weak-to-strong generalization (E.3), statistical validation (E.4), and case study (F).
4 Related Work
Multimodal Reasoning   MLLMs have demonstrated robust capabilities across diverse domains [58, 12, 38]. Despite these achievements, complex multimodal reasoning remains challenging due to its demands on both visual perception and high-level cognition. While recent approaches [10, 27] implement structured reasoning to enhance CoT capabilities [4], they face inherent limitations: (i) explicit search methods [14, 15] that suffer from computational inefficiency, and (ii) post-training methods [16, 18] that require substantial data and computational resources. These methods either use SFT which shows limited generalization, or employ RLVR like GRPO that lacks the integration of external knowledge [22]. Our work leverages high-level thought cards to seamlessly combine internal model capabilities with external reasoning guidelines. Moreover, many existing approaches employ rigid structures that limit flexibility across different tasks, overlooking the importance of adaptive reasoning in unleashing multimodal reasoning potential [65]. In contrast, our thought cards enable task-specific reasoning paths and adaptive selection, balancing performance with efficiency.
Tree-based Search   Tree structures have shown significant potential in language models [66, 67]. Recent research applies these tree-search methods to identify effective reasoning paths for MLLMs. While AR-MCTS [14] enhances multimodal reasoning by integrating MCTS with active retrieval, its extensive iteration and computational demands limit practical applications. Similarly, Mulberry [15] employs tree structures to distill 260K reasoning chains from models like GPT-4o, but requires substantial computational resources and high-capacity teacher models. These methods fail to balance performance with efficiency. To address these limitations, we propose incorporating high-level problem-solving guidelines into MCTS, achieving competitive performance with higher efficiency. While MCTS serves as our reasoning organization structure in this work, our framework is flexible enough to accommodate other advanced structures in future explorations.